### 08 Dec 2023 v1

# The defect of a cubic threefold

Abstract.
The defect of a cubic threefold $X$ with isolated singularities is a global invariant that measures the failure of $\mathbb{Q}$-factoriality. We compute the defect for such cubics by projecting from a singularity. We determine the mixed Hodge structure on the middle cohomology of $X$ in terms of both the defect and local invariants of the singularities. We then relate the defect to various geometric properties of $X$: in particular, we show that a cubic threefold is not $\mathbb{Q}$-factorial if and only if it contains either a plane or a cubic scroll.
1. Introduction
Cubic hypersurfaces have a distinguished place in the study of algebraic geometry. In particular, they are closely related to hyperkähler varieties. For a smooth cubic fourfold $V\subset\mathbb{P}^{5}$, one can consider an associated compactified intermediate Jacobian fibration $\pi:\mathcal{J}_{V}\rightarrow(\mathbb{P}^{5})^{*}$, constructed in [LSV17] and [Sac23]. When $V$ is general, this Lagrangian fibration has irreducible fibers, however this is no longer true without the generality assumption. In [Bro16], Brosnan gives a necessary condition for a fiber to be reducible: the cubic threefold obtained as a hyperplane section $X=H\cap V$ has positive defect, $\sigma(X):=b_{4}(X)-b_{2}(X)>0.$
Inspired by this observation, we study the defect of singular cubic threefolds $X\subset\mathbb{P}^{4}.$ We consider the case when $X$ has isolated (and thus rational) singularities; this is equivalent to $X$ appearing as a hyperplane section of a smooth cubic fourfold [Bro16, Lemma 4.7]. The possible combinations of isolated singularities that can occur were classified in [Vik23] (see also [Yok02], [All03] for a GIT perspective). The defect $\sigma(X)$ is a global invariant of $X$, depending on both the singularities and their position; and can also be defined as
$\sigma(X)=\mathrm{rank}(\mathrm{Weil}(X)/\mathrm{Cart}(X)).$
If the defect is zero, $X$ is $\mathbb{Q}$-factorial [Che09, Che10]; for a cubic threefold this is equivalent to being factorial (see for instance [Kaw88]). The geometry of $\mathbb{Q}$-factorial Fano threefolds has been well studied (see [Shr08], [CP10], [Kal11], [CPS19]). A cubic threefold with only ordinary double points in general position is $\mathbb{Q}$-factorial if and only if the number of points is less than 4 [Che10, Theorem 1.4]; for other combinations of singularities, the situation is more complex.
The defect of Fano threefolds and of hypersurfaces (in particular Calabi-Yau threefolds) has been studied in relation to their mixed Hodge structures ([Cle83], [Dim90], [NS95], [Cyn01]). The first result of this paper is the computation of the mixed Hodge structure on the middle cohomology of a cubic threefold $X\subset\mathbb{P}^{4}$ in terms of the singularity type of $X$, along with the defect.
Theorem 1.1.
Let $X$ be a cubic threefold with only isolated singularities. Then the Hodge-du Bois numbers of the mixed Hodge Structure on $H^{3}(X)$ are completely determined by a global invariant $\sigma(X),$ and two local invariants, namely the Du Bois and link invariants of the singularities.
Further, let $q\in\mathrm{Sing}(X)$ and denote by $k$ the number of irreducible components of the curve $C_{q}$ parametrising lines through $q$. Then:
$\sigma(X)=\begin{cases}k-1&\text{ if }\mathrm{corank}(q)\neq 2;\\ k-2&\text{ if }\mathrm{corank}(q)=2.\end{cases}$
This allows one to compute $\sigma(X)$ and the cohomology of $X$ in terms of the singularity type of $X$, along with the position. As a by-product, we obtain a classification of $\mathbb{Q}$-factorial cubic threefolds.
Our next goal is to determine when a cubic fourfold $V\subset\mathbb{P}^{5}$ contains a hyperplane section $X=V\cap H$ with $\sigma(X)>0$. It is clear that such a hyperplane section exists if $V$ contains a plane or a rational normal cubic scroll, as observed in [MO22], [Sac23]. We prove that this is an equivalence:
Theorem 1.2.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold with isolated singularities. Then the following are equivalent:
  1. (1)
    $\sigma(X)>0$;
  2. (2)
    $X$ contains either a plane, or a rational normal cubic scroll.
Remark 1.3.
It is possible for a cubic threefold with isolated singularities to contain surfaces of higher degree; indeed by [Aur90], [Koe92] such a cubic can contain an elliptic quintic scroll. This is realised by the Segre cubic for instance, see [Aur90, Remark 1.5] for more details.
Remark 1.4.
It was shown in [Mar23] that cubic fourfolds with involutions contain either many planes, or many cubic scrolls. It follows that their associated compactified intermediate fibrations have many reducible fibers, and the induced symplectic involutions (see [MM23a, MM23b]) are strictly birational.
In [CML09], the authors introduced the notion of a good line in a cubic threefold, in order to study degenerations of the intermediate Jacobian as $X$ becomes mildly singular (see also [CMGHL21]). The main tool is Mumford’s description of the intermediate Jacobian as a Prym variety: for a line on the cubic, projection exhibits the cubic as a conic bundle over $\mathbb{P}^{2}$, whose discriminant curve is a plane quintic with an associated double cover. For the choice of a good line, the singularities of the quintic are in one-to-one correspondence with those of $X$, and one can study the degeneration of the intermediate Jacobian as a Prym variety of the associated double cover, even in the singular situation [Bea77]. This notion was also used in [LSV17], who introduced the stronger notion of a very good line: for such a line, the associated quintic double cover is étale and has both curves irreducible. We show that such a line exists on a cubic threefold $X$, making the Prym construction possible, only when $\sigma(X)=0$.
Theorem 1.5.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold with isolated singularities. Assume further that the singularities are of corank at most 2. Then the conditions of Theorem 1.2 are equivalent to:
  1. (3)
    $X$ does not have a very good line.
This sharpens the results of [CML09] and [LSV17]: there are examples of cubic threefolds with worse than allowable (i.e not semi-stable) that none the less admit a very good line.
Outline
We recall relevant notation and definitions in §2. In §3.1 we prove our formula for the defect in terms of components of the associated curve $C_{q}$, and in §3.2 we provide some applications to $\mathbb{Q}$-factoriality of cubic threefolds. In §3.3 we compute the mixed Hodge structure for cubic threefolds, completing the proof of Theorem 1.1. We prove Theorem 1.2 in §4, and finally discuss the existence of very good lines in §5.
Acknowledgements
The authors would like to thank Radu Laza for valuable advice. They would like to thank Evgeny Shinder for useful discussions about $\mathbb{Q}$-factoriality and to thank both him and Ivan Cheltsov for helpful comments on a preliminary version. The second author is supported by Methusalem grant METH/21/03 – long term structural funding of the Flemish Government.
2. Preliminaries
Throughout, $X\subset\mathbb{P}^{4}$ is a cubic threefold with isolated singularities that is not a cone over a cubic surface. We recall the projection method in Section 2.1, which we will use to compute the cohomology of $X$. In Section 2.2 we define the local invariants attached to the singularities of $X$.
2.1. The projection method
Let $q\in X$ be an isolated singularity, and choose coordinates so that $q=[0:0:0:0:1]$. The equation of $X$ is thus
(2.1) $f(x_{0},\dots,x_{4})=x_{4}q_{2}(x_{0},\dots,x_{3})+g_{3}(x_{0},\dots,x_{3}),$
where $q_{2},g_{3}$ are homogeneous polynomials of degree 2 and 3 respectively. Projection from the point $q$ gives a birational map $X\dashrightarrow\mathbb{P}^{3}$; we have the diagram:
[…]
Define $Q_{q}:=V(f_{2}(x_{0},\dots,x_{3}))\subset\mathbb{P}^{3}$, and $S_{q}:=V(g_{3}(x_{0},\dots,x_{3}))$. We let $C_{q}:=Q_{q}\cap S_{q}$ be the $(2,3)$ complete intersection curve, that parametrises lines in $X$ passing through $q$.
Lemma 2.2.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold with only isolated singularities that is not a cone over a cubic surface. Assume $q\in X$ is an isolated singularity. Then
$Bl_{q}X\cong Bl_{C_{q}}\mathbb{P}^{3}.$
Proof.
This is Proposition 3.1 of [CML09]. For a more detailed proof, see [Hav16, Theorem 2.1.1]. ∎
Definition 2.3.
Let $q\in X$ be a singular point. We define the corank of $q$ to be equal to the corank of the quadric $Q_{q}.$
This definition agrees with the usual definition of corank of a critical point of a function in this case. Note that if $q\in X$ is a point of corank 3, then $Q_{q}$ is a double plane and the curve $C_{q}$ is non-reduced.
We will use the following well-known lemma:
Lemma 2.4.
The corank of $q$ is equal to one if and only if $q$ is of $A_{r}$ type with $r>1.$ The corank of $q$ is equal to zero if and only if $q$ is an $A_{1}$ singularity.
2.2. Local Invariants
Let $(X,x)$ be a germ of an isolated threefold singularity and $\psi:(Y,E)\rightarrow(X,x)$ its resolution with $E$ a normal crossing divisor. We fix a representative $\psi:Y\rightarrow X$ such that $X$ is a contractible Stein space. Since we will be considering singular cubic threefolds, whose isolated singularities are rational, we will use the properties of rational singularities to simplify the computation of local invariants where possible.
Definition 2.5.
[Ste97] We define Du Bois invariants for $p\geq 0,q>0$ as
$b^{p,q}(X,x):=\dim H^{q}(Y,\Omega^{p}_{Y}(\log E)(-E)).$
Recall that $(X,x)$ is a Du Bois singularity if $b^{0,q}=0$ for $30<q<$ 3, and $(X,x)$ is a rational singularity if $H^{i}(Y,\mathcal{O}_{Y})=0$ for $30<i<$ 3. Note that rational implies Du Bois.
Lemma 2.6.
Let $(X,x)$ be a rational singularity. Then $b^{p,q}=0$ except for possibly $b^{1,1}$ and $b^{2,1}$.
Proof.
By assumption, $b^{0,q}=0$ for $30<q<$ 3, and by [Ste97, Theorem 1] $b^{p,q}=0$ for $p+q>3$ or $q\geq 3$. We also have $b^{1,2}=0$ for a Du Bois singularity (see for instance [Ste97, p. 1369]). ∎
Definition 2.7.
Let $L$ denote the link of the singularity $(X,x)$. We define link invariants:
$l^{p,q}:=\dim H^{q}(E,\Omega^{p}(\log E)\otimes\mathcal{O}_{E})=\dim\mathrm{Gr}^{p}_{F}H^{p+q}(L),$
where $L$ is the link of the singularity.
Using Poincaré duality for $L$ we see that
$l^{p,q}=l^{3-p,2-q}.$
Definition 2.8.
The Deligne-Hodge numbers of the link $L$ are given by
$h_{i}^{p,q}(L)=\dim\mathrm{Gr}_{F}^{p}\mathrm{Gr}^{W}_{p+q}H^{i}(L).$
We see that $h_{i}^{p,q}=h_{i}^{q,p},$ and $l^{p,i-p}=\sum_{q}h_{i}^{p,q}$. Further, by semi-purity we have $h_{i}^{p,q}=0$ if $i<3$ and $p+q>3$.
Definition 2.9.
The Hodge-du Bois numbers of a compact complex algebraic variety $X$ are defined by:
$\underline{h}^{p,q}(X):=\dim\mathrm{Gr}_{F}^{p}H^{p+q}(X).$
If $X$ is smooth they recover the usual Hodge numbers $\underline{h}^{p,q}(X)=h^{p,q}(X)$. We note:
$\underline{h}^{p,q}(L)=\sum_{0\leq r\leq q}h^{p,r}_{p+q}(L).$
Lemma 2.10.
Let $(X,x)$ be a rational isolated local complete intersection singularity. Then $l^{p,q}=0$ except possibly for $l^{1,1}$.
Proof.
The link of a $3$-dimensional isolated complete intersection singularity is simply connected, so $H^{i}(L)\neq 0$ only for $i\in\{0,2,3,5\}$. It follows in this case that $l^{p,q}=0$ unless $p+q\in\{0,2,3,5\}.$ Since $(X,x)$ is a rational singularity, by [Ste97, Lemma 2] $l^{0,i}=l^{i,0}=0$ for all $i$. ∎
Definition 2.11.
For $(X,x)$ an isolated local complete intersection singularity, we let $M_{x}$ denote the Milnor fiber of the singularity. Then $H^{3}(M_{x})$ admits a mixed Hodge structure, and we set
$s_{p}:=\dim\mathrm{Gr}^{p}_{F}H^{n}(M_{x}).$
In the case of interest, i.e. cubic threefolds with isolated singularities, we can use the following theorem to relate these invariants together.
Theorem 2.12.
[Ste97, Theorem 6] Let $(X,x)$ be a three-dimensional Du Bois isolated complete intersection singularity with $l^{2,0}=0$. Then the Hodge numbers of the Milnor fiber are given by:
$s_{0}=0,\,\,s_{1}=b^{1,1},\,\,s_{2}=b^{1,1}+l^{1,1},\,\,s_{3}=l^{0,2}.$
3. The Cohomology of $X$
We compute the cohomology for a cubic threefold $X\subset\mathbb{P}^{4}$ with isolated singularities (that is not a cone over a cubic surface). First we define the defect $\sigma(X)$ and provide in §3.2 a simple way to compute this invariant by counting irreducible components of the $(2,3)$ complete intersection curve $C_{q}$ associated to a singular point $q\in X.$ In §3.3, we compute the mixed Hodge structure on $H^{3}(X,\mathbb{Z})$ in terms of local invariants of the singularities, as well as $\sigma(X).$ Further, we show that the mixed Hodge structure on $H^{3}(X,\mathbb{Z})$ is completely determined by that on $H^{1}(C_{q})$.
3.1. The defect of $X$
Definition 3.1.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold. The defect $\sigma(X)$ of $X$ is the rank of the finitely generated abelian group $\mathrm{Weil}(X)/\mathrm{Cart}(X)$.
By [NS95, Theorem 3.2], one can calculate the defect by $\sigma(X)=b_{4}(X)-b_{2}(X).$ We provide a geometric way to compute the defect for a cubic threefold with isolated singularities.
Proposition 3.2.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold with only isolated singularities, that is not the cone over a cubic surface. Let $q\in X$ be such a singularity, and let $k$ denote the number of irreducible components of the associated $(2,3)$ complete intersection curve $C_{q}$. Then:
$\sigma(X)=\begin{cases}k-1&\text{ if }\mathrm{corank}(q)\neq 2;\\ k-2&\text{ if }\mathrm{corank}(q)=2.\end{cases}$
Proof.
By Lemma 2.2, $\widetilde{X}:=Bl_{q}X\cong Bl_{C_{q}}\mathbb{P}^{3}$, and so we have two discriminant squares:
[…]
$H^{3}(E,\mathbb{Z})\cong h\cdot f^{*}H^{1}(C,\mathbb{Z}),$
via pullback and the cup-product, both of which are morphisms of mixed Hodge structures. Now we apply the Mayer-Vietoris sequence (see [PS08, Corollary-Definition 5.37]) to $(b)$:
$0\rightarrow H^{3}(\widetilde{X})\rightarrow H^{3}(E)\rightarrow H^{4}(\mathbb{P}^{3})\rightarrow H^{4}(\widetilde{X})\rightarrow H^{4}(E)\rightarrow 0$
By the above discussion, we conclude that $h^{3}(E)=h^{1}(C)$ and $h^{4}(E)=h^{2}(C)$. Since $H^{4}(\mathbb{P}^{3})$ has a pure Hodge structure of weight $4,$ and $H^{3}(E)$ has weights at most $3$, it follows that
$h^{3}(\widetilde{X})$ $=h^{3}(E)=h^{1}(C),$
$h^{4}(\widetilde{X})$ $=h^{2}(C)+1.$
We apply the same sequence to $(a)$; since $H^{3}(Q)=0$ this reduces to the short exact sequence
$0\rightarrow H^{4}(X)\rightarrow H^{4}(\widetilde{X})\rightarrow H^{4}(Q)\rightarrow 0$
Combining these facts we see that $h^{4}(X)=k+1-h^{4}(Q)$; since $h^{2}(X)=1$, the claim follows. ∎
3.2. $\mathbb{Q}$-factoriality
As an application of Proposition 3.2, and utilising the classification of possible singularities of a cubic threefold $X$, we are able to determine which cubic threefolds with isolated singularities are $\mathbb{Q}$-factorial. All of these results directly follow from studying the curve $C_{q}$ and counting the number of irreducible components. This is essentially contained in [Vik23, §2, 3 and 4]; we omit the details for brevity.
Proposition 3.3.
Let $X$ be a cubic threefold with isolated singularities, and assume that one of the singular points of $X$ has corank greater than 1.
  1. (1)
    If $X$ is a cone over a smooth cubic surface, then $\sigma(X)=6$. Further, $Weil(X)/Cart(X)$ is generated by the planes obtained as cones over the lines on the cubic surface.
  2. (2)
    If $q\in X$ is a corank 3 singularity, then $X$ is $\mathbb{Q}$-factorial if and only if $q$ is a $T_{333}$, $T_{344}$ or $Q_{10}$ singularity.
  3. (3)
    If $q\in X$ is a corank 2 non-$ADE$ singularity, then $\sigma(X)>0$.
Moreover, in all the cases above $\sigma(X)$ depends only on the singularity type of $X$, or in other words, $\sigma(X)$ is determined by the local data.
For combinations of $ADE$ singularities, $\sigma(X)$ depends on the singularity type and the geometry of $X$, but can still be determined by studying the curve $C_{q}$. We will focus on the $\mathbb{Q}$-factoriality of maximal combinations of singularities - a combination is maximal in a certain class of combinations of singularities if we cannot obtain this combination of singularities by deforming a cubic hypersurface with another combination of singularities from this class. One can easily extract the results for all combinations using [Vik23].
Remark 3.4.
Upper semicontinuity of cohomology implies that if a cubic threefold has defect 0, then all of its deformations have defect 0.
Proposition 3.5.
A cubic threefold with $E_{8}+A_{2}$ or $D_{5}+2A_{2}$ singularities is $\mathbb{Q}$-factorial. A cubic threefold with $E_{7}+2A_{1}$ is $\mathbb{Q}$-factorial if and only if it doesn’t contain a plane. Moreover, these are the maximal combinations containing a $D_{n}$ or $E_{n}$ singularity that can be $\mathbb{Q}$-factorial.
We illustrate the dependence on the geometry of $X$ with the example below:
Example 3.6.
A cubic threefold $X$ with a $D_{4}+2A_{1}$ combination of singularities can have $\sigma(X)=0$ or $1$. Explicitly, let $q\in X$ be of type $D_{4}$. By [Vik23, Section 3.2], we have that $Q_{q}=P_{1}\cup P_{2}$ for planes $P_{i}$ intersecting in a line $L$. The curve $C_{q}$ is a union of two plane cubics $C_{1}$ and $C_{2}$ which intersect at three distinct points on $L$. If $C_{i}$ are both irreducible with one node, then $\sigma(X)=0$. If $C_{1}$ is smooth and $C_{2}$ has two nodes (and thus reducible), then $\sigma(X)=1$. Both situations occur, for explicit cubics below:
$\sigma(X)$ $=0:\,\,x_{0}x_{1}x_{2}+x_{1}^{2}x_{4}-2x_{1}x_{4}^{2}+x_{2}^{2}x_{4}-2x_{2}x_{4}^{2}-x_{3}^{3}-x_{3}^{2}x_{4}+x_{4}^{3}=0;$
$\sigma(X)$ $=1:\,\,x_{0}x_{1}x_{2}+x_{1}x_{3}x_{4}+x_{2}^{3}-x_{3}^{3}+x_{3}x_{4}^{2}=0.$
Proposition 3.7.
A cubic threefold with $A_{10}$, $A_{8}+A_{2}$, $A_{6}+A_{4}$, $2A_{4}+A_{2}$ or $5A_{2}$ singularities is $\mathbb{Q}$-factorial. Moreover, these are the maximal constellations of $A_{n}$ singularities that can be $\mathbb{Q}$-factorial.
Remark 3.8.
In [KPS21]), the authors relate the property of $\sigma(X)=0$ to the existence of certain semiorthogonal decompositions of the bounded derived category of $X$, namely Kawamata type decompositions [KPS21, Theorem 1.1]. In particular, if $X$ is a cubic threefold with such a decomposition then $X$ has enough Weil divisors. One obtains more examples of cubic threefolds with singularities worse than nodes that do not admit such decompositions: for example, there exist a $\mathbb{Q}$-factorial cubic threefold with an $A_{9}$ singularity.
3.3. Mixed Hodge Structure on $H^{3}(X,\mathbb{Z})$
To completely understand the cohomology of the cubic threefold $X$ with isolated singularities, it remains to understand the mixed Hodge structure on the middle cohomology. Let $\{p_{1},\dots p_{r}\}\in X$ be the isolated singular points of $X$. We denote by
$B:=\sum_{i=1}^{r}b^{1,1}(X,p_{i})\text{ and }L:=\sum_{i=1}^{r}l^{1,1}(M_{p_{i}})$
the total Du Bois and link invariants of $X$.
Theorem 3.9.
Let $X$ be a cubic threefold with isolated singularities which is not a cone over a cubic surface. Then the non-zero Hodge-du Bois numbers of the mixed Hodge structure on $H^{3}(X)$ are computed as follows:
$\underline{h}^{2,1}(X)$ $=5-B;$
$\underline{h}^{1,2}(X)$ $=5-(L-\sigma)-B.$
Proof.
Recall that $X$ has rational hypersurface singularities. Since $X$ is a compact variety, it follows that $\mathrm{Gr}^{W}_{k}H^{3}(Z)=0$ for $k>3$. We let $\mathcal{X}\rightarrow\Delta$ be a $1$-parameter smoothing of $X$, and consider the vanishing cycle sequence:
$0\rightarrow H^{3}(X)\rightarrow H^{3}_{\lim}(X_{t})\rightarrow H^{3}_{\mathrm{van}}\rightarrow H^{4}(X)\rightarrow H^{4}_{\lim}(X_{t})\rightarrow 0$
Let us first consider the limit cohomology $H^{3}_{\lim}(X_{t})$. For fixed $p$, we have that
$\sum_{q}h^{p,q}(H^{3}_{\lim}(X_{t}))=h^{p,3-p}(X_{\infty}),$
where $X_{\infty}$ is a smooth cubic threefold. Denote by: $\alpha:=h^{1,1}(H^{3}_{\lim}(X_{t}))\text{ and }\beta:=h^{1,2}(H^{3}_{\lim}(X_{t}));$ it follows that $\alpha+\beta=5$.
The vanishing cohomology $H^{3}_{van}$ has a mixed Hodge structure induced through the isomorphism $H^{3}_{van}\cong\bigoplus H^{3}(M_{p_{i}})$, and we can compute the Hodge numbers using Theorem 2.12. In particular, $H^{3}_{\mathrm{van}}$ has no weight 2 cohomology. It follows that $h^{1,1}(H^{3}(X))=\alpha.$
On the other hand, the map $H^{3}_{van}\rightarrow H^{4}(X)$ is a morphism of pure weight 4 Hodge structures by [KLS21, Prop 5.5]. Since $\dim\ker(H^{4}(X)\rightarrow H^{4}_{\lim}(X_{t}))=\sigma(X)$ and is pure of type $(2,2)$ we have that $\alpha=L-\sigma$. Comparing weight three cohomology gives $\beta=h^{3,1}(H^{3}(X))+B$, and the result follows. ∎
The topology of $C_{q}$ determines the global invariant $\sigma(X)$; in fact, the local invariants $B,L$ are also determined by the singularities of $C_{q}$. More precisely:
Proposition 3.10.
There is an isomorphism $\mathrm{Gr}_{3}^{W}H^{3}(X)\cong\mathrm{Gr}_{1}^{W}H^{1}(C_{q})(-1)$ of Hodge Structures.
Proof of Proposition 3.10.
Let us apply the Mayer-Vietoris sequence to the discriminant square $(a)$ of Proposition 3.2:
[…]
Since $X$ is a hypersurface with isolated singularities, $H^{2}(X)$ has a pure weight 2 Hodge structure [KL20, Lemma 1.5], and $h^{2}(X)=1$. Since $\widetilde{X}=Bl_{q}X$, it follows that $h^{2}(\widetilde{X})=2$. We see that the sequence $H^{2}(X)\rightarrow H^{2}(\widetilde{X})\rightarrow H^{2}(Q)$ is an exact sequence of pure weight 2 Hodge structures; it follows that we have an isomorphism of mixed Hodge Structures: $H^{3}(X)\cong H^{3}(\widetilde{X}).$ Now let us apply the Mayer-Vietoris sequence to $(b)$:
[…]
Notice that $H^{4}(\mathbb{P}^{4})$ is a pure Hodge structure of weight $4$; since $H^{3}(E)$ has weights at most 3, it follows that we have an isomorphism of mixed Hodge Structures: $H^{3}(\widetilde{X})\cong H^{3}(E).$ In particular, $Gr_{3}^{W}H^{3}(X)\cong Gr_{3}^{W}H^{3}(E).$ We saw in the proof of Proposition 3.2 that $H^{3}(E,\mathbb{Z})\cong h\cdot f^{*}H^{1}(C_{q},\mathbb{Z}),$ via pullback and the cup-product, both of which are morphisms of mixed Hodge structures. In particular, it follows that we have an isomorphism of Hodge structures $Gr_{3}^{W}H^{3}(E)\cong Gr_{1}^{W}H^{1}(C_{q})(-1).$
4. Planes and rational normal cubic scrolls
Throughout, let $X$ be a cubic threefold with isolated singularities that is not a cone over a cubic surface. We will prove the following theorem:
Theorem 4.1 (Theorem 1.2).
Let $X$ be a cubic threefold with isolated singularities. Then the defect $\sigma(X)>0$ if and only if $X$ contains a plane or a rational normal cubic scroll.
Recall that a rational normal cubic scroll is a a smooth connected surface $\Sigma\subset\mathbb{P}^{4}$ of minimal degree with $\text{span}(\Sigma)=\mathbb{P}^{4}$. Such a surface is isomorphic to $\mathbb{P}^{2}$ blown up in a point. It is clear that if $X$ contains a plane or a cubic scroll then $\sigma(X)>0$ we include the proof for completeness.
Lemma 4.2.
If a cubic threefold $X$ contains a plane or a cubic scroll, then the defect $\sigma(X)>0.$
Proof.
Suppose that $\sigma(X)=0$; then every Weil divisor on $X$ is Cartier, and in particular for every Cartier divisor $Z\subset X$ there is a hypersurface $H\subset\mathbb{P}^{4}$ with $Z=X\cap H$ (see for [Che10, Remark 1.2] for instance). This is a contradiction: indeed, a plane can never be obtained as such an intersection, and a rational normal scroll is a nondegenerate surface of degree $3$. ∎
To prove the converse we again will exploit the geometry of the associated curve $C_{q}$. Combining Lemmas 4.3, 4.4, 4.5 below completes the proof of Theorem 4.1.
Lemma 4.3.
Let $X$ be a cubic threefold with only isolated nodal singularities, and suppose $\sigma(X)>0$. Then $X$ contains a plane or a rational normal cubic scroll.
Proof.
The case of nodes is essentially contained in [FW89]; we provide a proof using the language of this note for completion. Let $q\in X$ be a node, and project to obtain $C_{q}\subset\mathbb{P}^{3}$. The curve $C_{q}$ lies on a smooth quadric, and the singularities of $C_{q}$ are in one-to-one correspondence with the remaining singularities of $X$. Since $\sigma(X)>0$, the curve $C_{q}$ is reducible by Proposition 3.2. It follows that $C_{q}$ has at least $5$ nodes, thus $X$ must have at least $6$ nodes. If the number of nodes is strictly greater than $6$, the curve $C_{q}$ either contains a line, or a plane conic - both imply the existence of a plane in $X$. If $X$ has exactly $6$ nodes in general position, then $X$ contains a cubic scroll by [HT10, Prop. 23]. If the nodes are not in general position, then four nodes must be contained in a plane contained in $X$ by [Avi18, Lemma 2.10]. ∎
Lemma 4.4.
Let $X$ be a cubic threefold with a combination of $A_{n}$ singularities, with at least one with $n\geq 2$. Assume that $\sigma(X)>0$. Then $X$ contains a plane or a rational normal cubic scroll.
Proof.
Let $q\in X$ be the worst singularity (i.e $A_{n}$ for $n\geq 2)$; the curve $C_{q}\subset\mathbb{P}^{3}$ lies on the quadric cone $Q_{q}$. The blow up $\widetilde{Q}_{q}$ of $Q_{q}$ at the cone point is isomorphic to $\mathbb{F}_{2}$; let $\widetilde{C}\subset\mathbb{F}_{2}$ be the strict transform of $C_{q}\subset Q_{q}$. Applying Proposition 3.2, the curve $\widetilde{C}$ is reducible. Examining the proofs of [Vik23, Propositions 4.6–4.9], we see that whenever $\widetilde{C}$ is reducible, it contains either the premiage of a ruling $\gamma$ of $Q_{q},$ the preimage of a hyperplane section $\Gamma$ of $Q_{q}$ missing the cone point, or the preimage of two twisted cubics $\Gamma_{1},\Gamma_{2}\subset\mathbb{P}^{3}$.
In the first two cases, $X$ clearly contains a plane; it sufices to consider the latter. We claim that $X$ contains a rational normal cubic scroll. Indeed, $\Gamma_{1},\Gamma_{2}$ are two twisted cubics lying on the cubic surface $S_{q}$. Following the argument of [Dol16, Thm 3.5], the linear systems $|\Gamma_{1}|,|\Gamma_{2}|$ define two determinantal representations of the surface $S_{q}$, each obtained from the other by taking the transpose of the matrix. In an identical way to [Dol16, Thm 3.5], the cubic threefold $X$ is isomorphic to a hypersurface $V(\det(A))$, where $A$ is a $3\times 3$ matrix with linear forms in coordinates on $\mathbb{P}^{4}$. By [HT10, Prop 17], it follows that $X$ contains a rational normal cubic scroll (see also [Has00, Lemma 4.1.1]). ∎
Lemma 4.5.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold with at least one singularity $q\in X$ of corank $\geq 2$. Assume that $\sigma(X)>0.$ Then $X$ contains a plane.
Proof.
We project from $q$ the quadric surface $Q_{q}$ is either the union of two planes $P_{1}\cup P_{2}\subset\mathbb{P}^{3}$, or a double plane. In the first case, $C_{q}=C_{1}\cup C_{2}$, where $C_{i}\subset P_{1}$ are plane cubics, whereas in the second $C_{q}$ is a double cubic. Since $X$ has positive defect, in both cases it follows that $C_{q}$ contains a line component $L$, and $X$ contains a plane spanned by the lines through $q$ parametrised by $L$. ∎
5. Very good lines
In this section we relate Theorem 1.2 to another geometric property of the cubic threefold $X$, namely the non-existence of a very good line (see Definition 5.5). In particular, we will prove:
Theorem 5.1.
Let $X$ be a cubic threefold with isolated singularities. Assume $X$ does not have a singularity of type $T_{333},T_{334},Q_{10}$. Then $X$ has a very good line if and only if $\sigma(X)=0.$
This proves Theorem 1.5: the excluded singularities are those of corank $3$ for which $\sigma(X)=0.$
5.1. Very Good Lines
We begin by recalling the definitions.
Definition 5.2.
[CML09, Def 3.4],[LSV17, Def 2.2] A line $l\subset X$ is good if for any plane $P\subset\mathbb{P}^{4}$ containing $l$ and another line, the intersection $P\cap Y$ consists of three distinct lines.
Projecting from a line $l\subset X$ gives a conic bundle structure $f:Bl_{l}X\rightarrow\mathbb{P}^{2}$, where the discriminant curve $D_{l}\subset\mathbb{P}^{2}$ has degree $5$. This curve has a natural double cover $\widetilde{D}_{l}\rightarrow D$, where $\widetilde{D}_{l}$ parametrises the lines intersecting $l$.
Proposition 5.3.
[CML09, Prop 3.6] Let $l\subset X$ be a good line, $D_{l}$ and $\widetilde{D}_{l}$ as above. Then:
  1. (1)
    There exists a natural 1-to-1 correspondence between the singularities of $D_{l}$ and those of $X$, including the analytic type.
  2. (2)
    The double cover $\widetilde{D}_{l}\rightarrow D_{l}$ is étale.
Remark 5.4.
A cubic threefold with a singularity of corank $3$ does not have a good line - indeed if it did, then the quintic curve $D_{l}$ would have a singularity of corank 3, which is impossible.
Definition 5.5.
[LSV17, Defn 2.9] Let $X\subset\mathbb{P}^{4}$ be a cubic threefold, and $l\subset X$ a good line. Then $l$ is very good if the curve $\widetilde{D}_{l}$ of lines in $X$ meeting $l$ is irreducible.
We first establish the existence of good lines for a cubic threefold with $\sigma(X)=0$. First, we will use the following lemma, which seems to be well known but we could not locate a reference.
Lemma 5.6.
Let $X$ be a cubic threefold with $\sigma(X)=0,$ and assume there are no singularities of corank 3. Then $F(X)$ is irreducible.
Proof.
Assume first that at least one singular point $q\in X$ is of corank $1$, and let $C_{q}$ be the corresponding curve of lines through $q$. Since $\sigma(X)=0,$ by Proposition 3.2 the curve $C_{q}$ is irreducible. We claim $F(X)$ is birational to $\mathrm{Sym}^{2}(C_{q})$, and is thus irreducible. Let $l\in F(X)\setminus\mathrm{Sing}(F(X))$, and consider the plane $H:=\langle q,l\rangle\subset\mathbb{P}^{4}$. Note that because $\sigma(X)=0,$ this plane is not contained in $X$, and we intersect $H\cap X$. Since $X$ is singular at $q$, we have $H\cap X=l\cup l_{x}\cup l_{y}$, and $l_{x},l_{y}$ both contain $q$. Thus under the projection $\pi:Bl_{q}X\rightarrow\mathbb{P}^{3}$, both $\pi(l_{x})=x$ and $\pi(l_{y})$ are contained in $C_{q}$. Thus $\phi([l])=(x,y)$ defines a rational map $\phi:F(X)\dashrightarrow\mathrm{Sym}^{2}(C_{q}).$ We construct the rational inverse as follows: let $(x,y)\in\mathrm{Sym}^{2}(C_{q})$ where $x\neq y$. Then there exists lines $l_{x},l_{y}$ parametrised by $x,y\in C_{q}$ passing through the singular point $q;$ we take the residual line of the intersection $\langle l_{x},l_{y}\rangle\cap X.$
Now suppose that all singular points are corank $2$, and project from such a singular point $q\in X$. In this case, the quadric $Q_{q}=P_{1}\cup P_{2}$ where $P_{i}\cong\mathbb{P}^{2}$. Since $\sigma(X)=0$, by Proposition 3.2 the curve $C_{q}=C_{1}\cup C_{2}$ where $C_{i}\subset P_{i}$ is an irreducible cubic curve.
We claim that $F(X)$ is birational to $C_{1}\times C_{2},$ and is thus irreducible. Let $[l]\in F(X)\setminus\mathrm{Sing}(F(X)),$ and consider the plane $H:=\langle q,l\rangle\subset\mathbb{P}^{4}$. Again, intersecting $H\cap X=l\cup l_{x}\cup l_{y}$ where $l_{x},l_{y}$ are lines in $X$ passing through $q,$, and are thus parametrised by points $x,y\in C_{q}.$ Note that $\pi(l)$ is a line in $\mathbb{P}^{3}$, and thus intersects each $P_{i}$ in a single point ($\pi(l)$ cannot be contained in $P_{i}$ if it was, by analysing the equation of $X$ it must be a component of the $(2,3)$ intersection curve $C_{q}$).
It follows that exactly one of the points $x,y$ belong to each component $C_{i};$ we assume $x\in C_{1},$ $y\in C_{2}.$ Thus $\phi([l])=(x,y)$ defines a rational map
$\phi:F(X)\dashrightarrow C_{1}\times C_{2}.$
We construct the rational inverse: let $(x,y)\in C_{1}\times C_{2}$, with $x\neq y$. Then there exists lines $l_{x},l_{y}$ parametrised by $x,y\in C_{1}\cup C_{2}$ passing through the singular point $q$; we take the residual line of the intersection $\langle l_{x},l_{y}\rangle\cap X.$
Remark 5.7.
The Fano variety of lines has been well studied in the case of $\sigma(X)>0;$ in particular, [HT10] describes the components of $F(X)$ explicitly for the cubic threefold containing a rational normal cubic scroll (see also [Dol16]); in particular there are 3 components. The birational geometry of a Fano variety for a cubic threefold containing a plane appears in [Had00, §3.3], and in detail in [Bro23] over fields of characteristic not 2. Again, there are three components.
It follows that the general line on a cubic threefold with $\sigma(X)=0$ is a good line. More precisely:
Lemma 5.8.
Let $X$ be a cubic threefold with $\sigma(X)=0$, and assume there are no singularities of corank $3$. Then the set of good lines $U\subset F(X)$ is a (Zariski) open subset of $F(X).$
Proof.
Since $\sigma(X)=0,$ the Fano variety of lines $F(X)$ is reduced and irreducible by Lemma 5.6. The argument follows as in [CML09, Lemma 3.9]. ∎
5.2. Containing a plane or a cubic scroll
We first show that a cubic threefold containing a plane or a rational normal cubic scroll does not have a very good line.
Proposition 5.9.
Suppose that $X$ contains a plane or a rational normal cubic scroll. Then for every good line $l\in X$, the corresponding curve $\widetilde{D}_{l}$ is reducible.
Proof.
Since $X$ contains a plane or a cubic scroll, the Fano variety of lines is reducible (see Remark 5.7). Let $F(X):=\cup F_{i}$, where each $F_{i}$ is an irreducible surface. Recall that $F(X)$ is singular at $[m]\in F(X)$ if and only if the line $m\subset X$ passes through a singular point ([AK77, Cor 1.11] and[BVdV79, Cor 4]). First we will show that $\widetilde{D}_{l}$ is not contained in $\cap F_{i}\subset\mathrm{Sing}(F(X))$. Recall that by definition,
$\widetilde{D}_{l}:=\{[l^{\prime}]\in F(X)|l^{\prime}\cap l\neq\varnothing\}.$
Suppose $\widetilde{D}_{l}\subset\mathrm{Sing}(F(X));$ then every line $[l^{\prime}]\in\widetilde{D}_{l}$ passes through a singular point. Under the projection $f:Bl_{l}X\rightarrow\mathbb{P}^{2}$, the pairs of lines $l^{\prime},\tau(l^{\prime})$ (where here $\tau$ is the covering involution associated to $\widetilde{D}_{l}\rightarrow D_{l})$ map to a point on the discriminant curve $D_{l}$. Since $l$ is a good line, the singularities of $D_{l}$ are in one-to-one correspondence with that of $X$, hence the image under projection of a singular point $q$ is a singular point. Since every line in the fibers parametrised by $D_{l}$ passes through a singular point, every point of $D_{l}$ is singular - this is a contradiction due to $X$ having isolated singularities.
Assume now that $\widetilde{D}_{l}$ is irreducible. Then $\widetilde{D}_{l}$ is contained in a unique component of $F(X),$ say $F_{1}$. For $i\neq 1$, let
$Z_{i}:=\{x\in X:\exists[l^{\prime}]\in F_{i}\text{ with }x\in l^{\prime}\}.$
There are two possibilities, either $Z_{i}$ cover $X,$ or $Z_{i}$ is a surface contained in $X$. If $Z_{i}$ covers $X$, then for every point $x\in l$ there exists a line $[l^{\prime}]\in F_{i}$ intersecting $l$ at $x$ it follows that $\widetilde{D}_{l}$ has a component contained in $F_{i}$, and is thus reducible.
Now assume that $Z_{i}$ is a surface. Let $p:\mathbb{L}\rightarrow F(X)$ be the universal line over $F(X)$, with projection $q:\mathbb{L}\rightarrow X$. Note that $p$ is a $\mathbb{P}^{1}$ bundle over $F(X)$, and $q$ is generically finite. Let $p_{i}:\mathbb{L}_{i}\rightarrow F_{i}$, $q_{i}:\mathbb{L}_{i}\rightarrow X_{i}$ be the restrictions to the component $F_{i}$. Then $\mathbb{L}_{i}$ is a $\mathbb{P}^{1}$ bundle over the surface $F_{i}$, and $q_{i}$ is a dominant map to the surface $Z_{i}$. In particular, the generic fiber of $q_{i}$ is $1$-dimensional. Since $Z_{i}\subset X\subset\mathbb{P}^{4}$, the intersection of the good line $l$ and $Z_{i}$ is non empty; pick $z\in Z_{i}\cap l$, and consider the fiber $q_{i}^{-1}(z).$ Since the dimension of the fibers of $q_{i}$ is an upper-semi continuous function, $\dim q_{i}^{-1}(z)\geq 1.$ In particular, there is a $1-$parameter family of lines contained in $F_{i}$ that pass through $z$, parametrised by a curve $C_{z}\subset F_{i}$. Clearly, $C_{z}$ is a component of $\widetilde{D}_{l}$, and $\widetilde{D}_{l}$ is reducible. ∎
Using the equivalence of Theorem 1.2, this immediately gives the following corollary.
Corollary 5.10.
Let $X$ be a cubic fourfold with $\sigma(X)>0$. Then $X$ does not contain a very good line.
5.3. Assuming non-existence of a very good line
To complete the proof of Theorem 5.1, it remains to prove the converse: if $X$ does not have a very good line, then $\sigma(X)>0$. By [CML09, §3.2], $X$ has a very good line if $\mu(X)\leq 5$, or if the singularities of $X$ are allowable (see also [LSV17, §2]). Note that such cubic threefolds have $\sigma(X)=0$ by Proposition 3.2.
Definition 5.11.
A cubic threefold $X$ has allowable singularities if it has singularities of type $A_{k}$ with $k\leq 5$ or $D_{4}$ singularities.
We thus make the following assumption for the rest of this section:
Assumption 5.12.
The cubic threefold $X$ satisfies $\mu(X)\geq 6,$ there is at least one non-allowable singularity $q\in X$ and $X$ does not have a singularity with corank $3$.
Suppose that $l\subset X$ is a good line, but is not very good. Then by definition the associated curve $\widetilde{D}_{l}\subset F(X)$ has multiple components, but the discriminant curve $D_{l}:=\widetilde{D}_{l}/\tau\subset\mathbb{P}^{2}$ can be either reducible or irreducible. We consider each case separately.
Theorem 5.13.
Suppose that for all good lines $l\in X$, the discriminant curve $D_{l}$ is reducible. Then $X$ contains a plane or a cubic scroll.
In order to prove Theorem 5.13, there are two situations to consider:
  1. (1)
    for every good line $l$, the curve $D_{l}$ contains a line component $M_{l}\subset D_{l}$; or
  2. (2)
    for at least one good line the curve $D_{l}$ is the union of a smooth conic and an irreducible plane cubic.
Proposition 5.14.
Suppose that for every good line $l\subset X$, the quintic $D_{l}$ contains a line component $M_{l}.$ Then $\sigma(Y)>0.$
Proof.
Assume that $\sigma(X)=0.$ Then $F(X)$ is irreducible and reduced by Lemma 5.6, and the set of good lines is open in $F(X).$ For each good line $l\subset X$, let $D_{l}=M_{l}\cup\Gamma_{l}$, where $M_{l}$ is the line component, and $\Gamma_{l}$ is a (possibly reducible) quartic.
Denote by $S_{l}:=\pi^{-1}(M_{l})$ the preimage in $X$ of the line component under the projection map $\pi:Bl_{l}X\rightarrow\mathbb{P}^{2}$. Then $S_{l}$ is a cubic surface in $X$ that contains a 1-parameter family of (pairs) of lines that meet the given line $l$. If $S_{l}$ is reducible for some $l$, then $X$ contains a plane. Thus we assume $S_{l}$ is irreducible for every choice of good line $l$.
For every plane $P$ that contains $l$, the intersection $P\cap S_{l}$ has a residual singular conic. The singular point moves in $S$ – it follows by Bertini that $\mathrm{Sing}(S_{l})=M^{\prime}_{l}$ is a line. This line $M^{\prime}_{l}\subset S_{l}\subset X$ necessarily passes through a singular point of $X$; indeed, $M_{l}$ intersects the quartic $\Gamma_{l}$ in at least one point, whose preimage is a singular point of $X$ contained in $M^{\prime}_{l}.$ It follows that $X$ contains an infinite number of cubic surfaces that are singular along a line passing through a singular point of $X$. This is a contradiction by Lemma 5.15 below. ∎
Lemma 5.15.
Let $X$ be a cubic threefold with singularities of corank $\leq 2.$ Assume $\sigma(X)=0.$ Then $X$ contains only finitely many cubic surfaces which are singular along a line.
Proof.
This is essentially [LSV17, Lemma 2.6] – their proof only uses the fact that $F(X)$ is irreducible, and thus holds for $X$. The conclusion of the proof of the lemma is that this forces the curve $C_{q}$ to be nonreduced, which is true only in the corank 3 case. Thus the result holds. ∎
Proof of Theorem 5.13.
Assume that $X$ does not contain a plane or a cubic scroll. In particular, we assume $\sigma(X)=0.$ By Proposition 5.14, it suffices to consider the situation when there exists at least one good line $l\subset X$ such that $D_{l}$ is the union of an irreducible cubic $T_{l}$ and a smooth conic $Q_{l}$. We will analyse the singularities of $D_{l}$ and use Proposition 3.2 to conclude that the defect of $X$ should be positive, leading to a contradiction.
Since the component $T_{l}$ of $D_{l}$ is irreducible, it is either smooth or has one singularity. If $T_{l}$ is smooth, then $D_{l}$ has only $A_{n}$ singularities occuring as the six intersection points of $T_{l}$ and $Q_{l}$ (counted with multiplicities). The combination of $A_{n}$ singularities is determined by a partition of six into a sum: there are eleven possibilities. In the case of $6A_{1}$ singularities (i.e. the partition 1+1+1+1+1+1), $X$ contains a cubic scroll or a plane, since the singularities of $X$ are the same as those of $D_{l}$ by Proposition 5.3. In the remaining cases, the proofs of [Vik23, Propositions 4.6-4.9]) imply that the curve $C_{q}$ associated to any singularity $q$ of $X$ is reducible. By Proposition 3.2, it follows that $\sigma(X)>0,$ a contradiction.
Next we assume that $T_{l}$ has a node. If the smooth conic $Q_{l}$ does not pass through the node, then $X$ has a combination of $A_{n}$ singularities again, and we proceed as above. If $Q_{l}$ passes through the node, then the corresponding singularity of $D_{l}$ has type either $D_{4},D_{6},D_{8},D_{10}$ or $D_{12}$. Singularities of types $D_{n}$ with $n>8$ do not occur on cubic threefolds by [Vik23, Theorem I], and thus $D_{10}$ or $D_{12}$ do not occur on $D_{l}$. Analysing the possible combinations of singularities on $X$ following [Vik23, Section 3.2], we see immediately that $\sigma(X)>0$ in all the cases in question.
Finally, we can assume that $T_{l}$ has a cusp and $Q_{l}$ passes through the cusp. If the resulting singularity is $ADE$, then it is of $D_{5}$, $D_{7}$ or $E_{7}$ type. The case by case analysis again follows from [Vik23, Section 3.2] and leads to the conclusion that $\sigma(X)>0$. If the resulting singularity of $D_{l}$ is non-$ADE$ of corank 2, then projection from the corresponding singularity $q\in X$ results in a reducible curve $C_{q}$ with at least three components, also implying $\sigma(X)>0$. Corank 3 singularities are impossible since we are considering a plane curve. ∎
Theorem 5.1 follows from Theorem 5.13 and the following theorem:
Theorem 5.16.
Let $X$ be a cubic threefold with isolated singularities of corank $\leq 2.$ Suppose that for every good line $l\in X$ the curve $\widetilde{D}_{l}$ is reducible. Then $\sigma(X)>0$.
Proof.
It suffices to assume that $D_{l}\subset\mathbb{P}^{2}$ is an irreducible quintic plane curve for a general choice of good line $l\subset X$.
Let $l\subset X$ be a general good line. Since $\widetilde{D}_{l}\rightarrow D_{l}$ is étale double cover, $\widetilde{D}_{l}=\widetilde{D}_{l}^{1}\cup\widetilde{D}_{l}^{2}$. Assume for a contradiction that $\sigma(X)=0.$ This implies $F(X)$ is irreducible and reduced by Lemma 5.6. We follow the strategy of [LSV17, §2]: almost all of the arguments work with the assumption that $F(X)$ is irreducible and reduced.
Consider the incidence variety $P\subset F(X)\times X$, which is a $\mathbb{P}^{1}$ bundle $p:P\rightarrow F(X).$ Since $F(X)$ is irreducible and reduced, so is $P$. Note also that the degree of $q:P\rightarrow X$ is $6$ as for a smooth cubic threefold, and the degree of $pr_{2}:(P\times_{X}P)\setminus\Delta_{P}\rightarrow P$ is $5$. We first claim that the assumption of $\widetilde{D}_{l}$ being reducible implies that $(P\times_{X}P)\setminus\Delta_{X}$ has two irreducible components dominating $X$. Indeed, this is [LSV17, Lemma 2.12]; the two components are given as
$\mathcal{C}_{1}=\bigcup_{[l]\in F(X)}\widetilde{D}_{l}^{1}\text{ and }\mathcal{C}_{2}=\bigcup_{[l]\in F(X)}\widetilde{D}_{l}^{2}.$
Next, we will show that $(P\times_{X}P)\setminus\Delta_{P}$ is irreducible, giving a contradiction as in [LSV17, §2]. First, we note that every component $Z$ of $(P\times_{X}P)\setminus\Delta_{P}$ dominates $P$ by second projection. In [LSV17, Lemma 2.14], the authors claim that the only possibility for $Z$ not to dominate $P$ is if there is a curve $W\subset X$ such that for every $x\in W,$ there is a curve $C_{x}$ of lines passing through $x$. This is impossible by [LSV17, Claim 2.15]: there are only finitely many such points $x\in Y.$
Recall that we have the map
[…]
and the fiber over $[l]\in F(X)$ is identified with $\widetilde{D}_{l}.$ Consider the involution $i$ of $P\times_{X}P\setminus\Delta_{P}$ mapping $(l_{1},l_{2},x)$ where $l_{1}\cap l_{2}=x$ to $(l_{3},l_{2},z)$ where $l_{3}$ is the residual line of $\langle l_{1},l_{2}\rangle\cap X,$ and $z=l_{3}\cap l_{2}.$ When restricted to $\widetilde{D}_{l},$ this is the involution giving the double cover $\widetilde{D}_{l}\rightarrow D_{l}.$ By the proof of [LSV17, Lemma 2.16], since $D_{l}=\widetilde{D}_{l}/i$ is irreducible, then so is $(P\times_{X}P)\setminus\Delta_{P}/i$. It follows that $(P\times_{X}P)\setminus\Delta_{P}$ has at most two components that are exchanged under the involution.
We have shown that ($P\times_{X}P)\setminus\Delta_{P}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\bigcup_{[l]\in F(X)}\widetilde{D}_{l}^{1}$ and $\mathcal{C}_{2}=\bigcup_{[l]\in F(X)}\widetilde{D}_{l}^{2}$. Let $k_{1},k_{2}$ be the degrees of $pr_{2}:\mathcal{C}_{i}\rightarrow P$; we have either $(k_{1},k_{2})=(2,3),(1,4)$ since these are the only possible degrees of the components $(\widetilde{D}_{l}^{1},\widetilde{D}_{l}^{2})$ over $l$. To reach the contradiction, one applies the same arguments as in the proof of [LSV17, Lemma 2.13], which culminates in a contradiction of the possibility of the degrees $(k_{1},k_{2})$ – we omit the details.
Thus assuming $\widetilde{D}_{l}$ is reducible or all lines $l$, we have shown that if $\sigma(X)=0$ then $P\times_{X}P\setminus\Delta_{P}$ is both reducible and irreducible - a contradiction. Thus $\sigma(X)>0.$